OpenStudy (anonymous):

Use Green's Theorem to evaluate the line integral: int ( 2y dx + 5x dy ) where C is the circle (x-1)^2 + (y+3)^2 = 25

OpenStudy (anonymous):

\[\int\limits_{C}^{}2y dx + 5x dy\] around: \[(x-1)^2 + (y+3)^2 = 25\]

OpenStudy (anonymous):

I get to here: \[\int\limits_{}^{}\int\limits_{R}^{}3 dA\] but I'm not sure how to set the limits of this off-center circle.

OpenStudy (anonymous):

Nevermind, I've got it now.